# Learning Policy Committees for Effective Personalization in MDPs with Diverse Tasks

### v1

@@ Abstract
Many dynamic decision problems, such as robotic control, involve a series of tasks, many of which are unknown at training time. Typical approaches for these problems, such as multi-task and meta reinforcement learning, do not generalize well when the tasks are diverse. On the other hand, approaches that aim to tackle task diversity, such as using task embedding as policy context and task clustering, typically lack performance guarantees and require a large number of training tasks. To address these challenges, we propose a novel approach for learning a policy committee that includes at least one near-optimal policy with high probability for tasks encountered during execution. While we show that this problem is in general inapproximable, we present two practical algorithmic solutions. The first yields provable approximation and task sample complexity guarantees when tasks are low-dimensional (the best we can do due to inapproximability), whereas the second is a general and practical gradient-based approach. In addition, we provide a provable sample complexity bound for few-shot learning. Our experiments on MuJoCo and Meta-World show that the proposed approach outperforms state-of-the-art multi-task, meta-, and task clustering baselines in training, generalization, and few-shot learning, often by a large margin.
@@ 1 Introduction
Multi-Task RL (MTRL): The broad goal of MTRL approaches is to leverage inter-task relationships to learn policies that are effective across multiple tasks (Yang et al., 2020; Sodhani et al., 2021; Sun et al., 2022). An important challenge in MTRL is task interference. One class of approaches aims to mitigate this issue through gradient alignment techniques (Hessel et al., 2019; Yu et al., 2020a). An alternative series of approaches addresses this issue from a representation learning perspective, enabling the learning of policies that explicitly condition on task embeddings (Hendawy et al., 2024; Lan et al., 2024; Sodhani et al., 2021). However, most MTRL methods still face challenges when tasks are diverse, particularly when it comes to generalizing to previously unseen tasks.
@@
Reward-Free RL: Another related line of research is reward-free RL (Agarwal et al., 2023; Cheng et al., 2022; Jin et al., 2020). However, results in this space make strong assumptions, such as assuming that the action space is finite or linear dynamics. Our results, on the other hand, make no assumptions on the dynamics and do not depend on the size or dimension of the state or action space.
@@ 2 Model
where the expectation is with respect to $\mathcal{T}_{\tau}$ and $\rho$. Let $V^{*}_{\tau}$ denote an optimal policy for a task $\tau$. Define $V_{\tau}^{\Pi}=\max_{\pi\in\Pi}V_{\tau}^{\pi}$, that is, we let the value of a committee $\Pi$ to a task $\tau$ be the value of the best policy in the committee for this task. There are a number of reasons why this evaluation of a committee is reasonable. As an example, if a policy implements responses to prompts for conversational agents and $\Pi$ is small, we can present multiple responses if there is significant semantic disagreement among them, and let the user choose the most appropriate. In control settings, we can rely on domain experts who can use additional semantic information associated with each $\pi\in\Pi$ and the tasks, such as the descriptions of tasks $\pi$ was effective for at training time, and similar descriptions to test-time tasks, to choose a policy. Moreover, as we show in Section 3.4, this framework leads naturally to effective few-shot adaptation, which requires neither user nor expert input to determine the best policy.
@@
Unless two tasks $\tau_{1},\tau_{2}$ have only their reward functions $r_{1}$ and $r_{2}$ differ by potential shaping, positive linear scaling, and $S_{1}$-redistribution, we have $V^{\Pi}_{\tau_{1}}\geq V^{\pi}_{\tau_{1}}$, $V^{\Pi}_{\tau_{2}}\geq V^{\pi}_{\tau_{2}}$, and $V^{\Pi}_{\tau_{1}}+V^{\Pi}_{\tau_{1}}>V^{\pi}_{\tau_{1}}+V^{\pi}_{\tau_{2}}$ for any single policy $\pi$.
@@ 3.2 Clustering
$\min_{\begin{subarray}{c}\{\theta_{1},\ldots,\theta_{K}\};\\ \alpha\in\mathbb{R}^{nK}\end{subarray}}\sum_{i}\mathbf{ReLU}\left(\left\{\sum_{k=1}^{K}\sigma_{k}(\alpha_{i})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right),$ (2)
where $\sigma(\cdot)$ is a softmax function. Next, we demonstrate that this is a principled proxy by showing that when full coverage of $T$ is possible, solutions of (1) and (2) coincide. The proof is in Appendix E.
@@ 3.3 Theoretical Analysis for MTRL
Let $\pi_{i}^{*}$ denote the optimal policy for task $\pi_{i}$. We use $V_{i}^{\pi_{j}^{*}}$ to denote the value of task $\tau_{i}$ using a policy that is optimal for task $\tau_{j}$.
@@ 3.5 Training
The RL problem itself will typically dominate computational complexity in practice, but in cases where RL can be efficiently learned (Brafman & Tennenholtz, 2002; Kearns & Singh, 2002) and dimension $d$ is constant, we additional have polynomial algorithmic complexity.
@@ 4 Experiments
We study the effectiveness of our approach—pacman—in two environments, MuJoCo (Todorov et al., 2012) and Meta-World (pmlr-v100-yu20a). In the former, the tasks are low-dimensional and parametric, and we only vary the reward functions, whereas the latter has non-parametric robotic manipulation tasks with varying reward and transition dynamics.
@@ 4.3 Further Empirical Investigation of Our Algorithm
We investigate our algorithmic contribution in two ways. First, we compare our method for task clustering with three common clustering methods: KMeans++ (Arthur & Vassilvitskii, 2006), DBScan (Khan et al., 2014), GMM (Bishop, 2006), as well as with random clustering.
Table 3: Comparison of clustering algorithms in HalfCheetahVel.
Method 6 Million Frames 12 Million Frames
KMeans++ $-105.91\pm 3.44$ $-89.50\pm 3.10$
DBScan $-234.05\pm 9.56$ $-213.42\pm 4.09$
GMM $-239.32\pm 11.27$ $-199.86\pm 9.54$
Random $-274.13\pm 16.76$ $-258.07\pm 13.62$
PACMAN -97.42 $\pm$ 3.70 -74.20 $\pm$ 6.76
As shown in Table 3, the proposed approach outperforms all baselines. In Appendix H.2 we show that the distributional improvement is even more pronounced.
Additionally, we consider the impact of varying $K$, and show that efficacy of pacman is non-monotonic in $K$. The reason is that once $K$ is sufficiently large to cover the entire set of training tasks, increasing it further reduces the number of training tasks in individual clusters and thereby hurts generalization, both within clusters and to previously unseen tasks. Finally, we also consider the impact of changing the value of $\epsilon$, and observe that the results are relatively robust to small changes in $\epsilon$. See Appendix H for further details on these ablations.
@@ 5 Conclusion and Limitations
Theorem 2.6 from (Skalse & Abate, 2023) says that two tasks $\tau_{1},\tau_{2}$ have the same ordering of policies if and only if $r_{1}$ and $r_{2}$ differ by potential shaping, positive linear scaling, and $S^{\prime}$-redistribution. Therefore, if we could find optimal policies for these two tasks separately; they necessarily differ. And by forming a policy committee of these two optimal policies, we have $V^{\Pi}_{\tau_{1}}\geq V^{\pi}_{\tau_{1}}$, $V^{\Pi}_{\tau_{2}}\geq V^{\pi}_{\tau_{2}}$, and $V^{\Pi}_{\tau_{1}}+V^{\Pi}_{\tau_{1}}>V^{\pi}_{\tau_{1}}+V^{\pi}_{\tau_{2}}$ for any policy $\pi$. ∎
@@ Appendix C Pseudocode of Greedy Intersection Algorithm
We first show $A_{1}\subset A_{2}$. Pick any $\{\theta_{1},\dots,\theta_{K}\}\in A_{1}$. Due to the premise, for each $i$, since $\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}-\epsilon\leq 0$ , there exists $\theta_{k^{*}}$ such that $\|\theta_{k^{*}}-\theta_{i}\|_{\infty}-\epsilon\leq 0$. Thus, we can have $\alpha_{k^{*}}(i)=1$, and $\alpha_{k}(i)=0$ for all the other $k\neq k^{*}$. Then $\mathbf{ReLU}\left(\left\{\sum_{k\in[K]}\mathrm{softmax}_{k}(\alpha_{i})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right)=\mathbf{ReLU}(\|\theta_{k^{*}}-\theta_{i}\|_{\infty}-\epsilon)=0.$ By setting $\alpha$ this way, we could achieve the zero loss for the relaxation problem. Hence $\{\theta_{1},\dots,\theta_{K}\}\in A_{2}$.
Now to show $A_{2}\subset A_{1}$, suppose $\{\theta_{1},\dots,\theta_{K}\},\alpha$ is a optimal solution. Due to the premise, we must have that $\mathbf{ReLU}\left(\left\{\sum_{k\in[K]}\mathrm{softmax}_{k}(\alpha_{i})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right)=0$ for each $i$. Now fix $i$, since softmax($\alpha$) is nonegative and sums to $1$, there must be some positive coordinate $\mathrm{softmax}_{k^{\prime}}(\alpha_{i})$. Hence for all such $k^{\prime}$, $\mathbf{ReLU}(\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}-\epsilon)=0$, i.e., $\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}\leq\epsilon$. Thus, $\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}\leq\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}\leq\epsilon$ also holds, and $\{\theta_{1},\ldots,\theta_{K}\}\sum_{i}\mathbf{1}(\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}\leq\epsilon)=n$. Consequently, $\{\theta_{1},\dots,\theta_{K}\}\in A_{1}$. ∎
@@ H.2 Additional results for empirical investigation of our method
H.2.1 Ablations over clustering methods
@@
The advantage of our algorithm becomes even more apparent when focusing on the welfare of the majority. To illustrate this, we present a histogram of rewards for individual test tasks during zero-shot testing using policies trained with our algorithm versus KMeans++ on the Half-Cheetah (velocity) benchmark.
@@ H.2.2 Hyperparameter Ablations
Finally, we show the effect of the $\epsilon$ hyperparmater in the Meta-World zero shot setting. These results are reported for success rate across all tasks.
@@
Table 5: Performance metrics for different $\epsilon$ values at 500K and 1M steps.

### v2

@@ Abstract
Many dynamic decision problems, such as robotic control, involve a series of tasks, many of which are unknown at training time. Typical approaches for these problems, such as multi-task and meta-reinforcement learning, do not generalize well when the tasks are diverse. On the other hand, approaches that aim to tackle task diversity, such as using task embedding as policy context and task clustering, typically lack performance guarantees and require a large number of training tasks. To address these challenges, we propose a novel approach for learning a policy committee that includes at least one near-optimal policy with high probability for tasks encountered during execution. While we show that this problem is in general inapproximable, we present two practical algorithmic solutions. The first yields provable approximation and task sample complexity guarantees when tasks are low-dimensional (the best we can do due to inapproximability), whereas the second is a general and practical gradient-based approach. In addition, we provide a provable sample complexity bound for few-shot learning. Our experiments on MuJoCo and Meta-World show that the proposed approach outperforms state-of-the-art multi-task, meta-, and task clustering baselines in training, generalization, and few-shot learning, often by a large margin. Our code is available at https://github.com/CERL-WUSTL/PACMAN/.
@@ 1 Introduction
Multi-Task RL (MTRL): The broad goal of MTRL approaches is to leverage inter-task relationships to learn policies that are effective across multiple tasks (Yang et al., 2020; Sodhani et al., 2021; Sun et al., 2022). An important challenge in MTRL is task interference. One class of approaches aims to mitigate this issue through gradient alignment techniques (Hessel et al., 2019; Yu et al., 2020a). An alternative series of approaches addresses this issue from a representation learning perspective, enabling the learning of policies that explicitly condition on task embeddings (Hendawy et al., 2024; Lan et al., 2024; Sodhani et al., 2021).
However, most MTRL methods still face challenges when tasks are diverse, particularly when it comes to generalizing to previously unseen tasks.
@@
Reward-Free RL: Another related line of research is reward-free RL where the agent’s goal is to first explore the MDP without a pre-specified reward function and then plan near-optimal policies for a set of given reward functions (Agarwal et al., 2023; Cheng et al., 2022; Jin et al., 2020). However, results in this space make strong assumptions, such as assuming that the action space is finite or linear dynamics. Our results, on the other hand, make no assumptions on the dynamics and do not depend on the size or dimension of the state or action space.
@@ 2 Model
where the expectation is with respect to $\mathcal{T}_{\tau}$ and $\rho$. Let $\mathcal{P}$ be a (possibly restricted, e.g., parametric) space of policies that we take to be exogenously specified. We define $V^{*}_{\tau}=\max_{\pi\in\mathcal{P}}V_{\tau}^{\pi}$ as the optimal value for a task $\tau$, that is, the value of an optimal policy for $\tau$.
Define $V_{\tau}^{\Pi}=\max_{\pi\in\Pi}V_{\tau}^{\pi}$, that is, we let the value of a committee $\Pi$ to a task $\tau$ be the value of the best policy in the committee for this task. There are a number of reasons why this evaluation of a committee is reasonable. As an example, if a policy implements responses to prompts for conversational agents and $\Pi$ is small, we can present multiple responses if there is significant semantic disagreement among them, and let the user choose the most appropriate. In control settings, we can rely on domain experts who can use additional semantic information associated with each $\pi\in\Pi$ and the tasks, such as the descriptions of tasks $\pi$ was effective for at training time, and similar descriptions to test-time tasks, to choose a policy. Moreover, as we show in Section 3.4, this framework leads naturally to effective few-shot adaptation, which requires neither user nor expert input to determine the best policy.
@@
Suppose $K>1$. Unless two tasks $\tau_{1},\tau_{2}$ have only their reward functions $r_{1}$ and $r_{2}$ differ by potential shaping, positive linear scaling, and $S_{1}$-redistribution, we have $V^{\Pi}_{\tau_{1}}\geq V^{\pi}_{\tau_{1}}$, $V^{\Pi}_{\tau_{2}}\geq V^{\pi}_{\tau_{2}}$, and $V^{\Pi}_{\tau_{1}}+V^{\Pi}_{\tau_{2}}>V^{\pi}_{\tau_{1}}+V^{\pi}_{\tau_{2}}$ for any single policy $\pi$.
@@ 3.2 Clustering
$\min_{\begin{subarray}{c}\{\theta_{1},\ldots,\theta_{K}\};\\ w\in\mathbb{R}^{nK}\end{subarray}}\sum_{i}\mathbf{ReLU}\left(\left\{\sum_{k=1}^{K}\sigma(w_{ik})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right),$ (2)
where $w$ is the assignment matrix mapping each $\theta_{i}$ to $\theta_{k}$ with unbounded weights and $\sigma(\cdot)$ is a softmax function to normalize the assignment. Next, we demonstrate that this is a principled proxy by showing that when full coverage of $T$ is possible, solutions of (1) and (2) coincide. The proof is in Appendix E.
@@ 3.3 Theoretical Analysis for MTRL
Let $\pi_{i}^{*}$ denote the optimal policy for task $\tau_{i}$. We use $V_{i}^{\pi_{j}^{*}}$ to denote the value of task $\tau_{i}$ using a policy that is optimal for task $\tau_{j}$.
@@ 3.5 Training
Furthermore, we emphasize that the additional overhead introduced by our method (due to clustering) is small, with the computational complexity in practice being predominantly determined by the RL problem. To illustrate, our Meta-World experiments show that training a single policy for 1 million steps necessitates approximately 40 hours using an A40 GPU. Conversely, the clustering process completes in roughly 1 second within a Google Colab notebook, and obtaining task embeddings takes approximately 2 minutes on an A40 GPU. Importantly, in cases where efficient RL learning is achievable (Brafman & Tennenholtz, 2002; Kearns & Singh, 2002) and the dimension d remains constant, our approach additionally boasts polynomial algorithmic complexity.
@@ 4 Experiments
We study the effectiveness of our approach—pacman—in two environments, MuJoCo (Todorov et al., 2012) and Meta-World (Yu et al., 2020b). In the former, the tasks are low-dimensional and parametric, and we only vary the reward functions, whereas the latter has non-parametric robotic manipulation tasks with varying reward and transition dynamics.
@@
Additionally, we highlight that we use success rate instead of returns as our evaluation metric in Meta-World, as it has been the standard metric across different papers.
@@ 4.3 Further Empirical Investigation of Our Algorithm
We investigate our algorithmic contribution in two ways. First, we compare our method with three common clustering methods in Meta-World’s zero-shot setting: KMeans++ (Arthur & Vassilvitskii, 2006), DBScan (Khan et al., 2014), GMM (Bishop, 2006), as well as with random clustering.
Method 125K Steps 250K Steps
KMeans++ $0.22\pm 0.06$ $0.30\pm 0.07$
DBScan $0.19\pm 0.04$ $0.28\pm 0.06$
GMM $0.29\pm 0.07$ $0.33\pm 0.08$
Random $0.22\pm 0.04$ $0.25\pm 0.04$
PACMAN (Ours) 0.36 $\pm$ 0.07 0.48 $\pm$ 0.09
Table 3: Performance Comparison for different clustering methods at 125K and 250K Steps.
Notably, pacman exhibits $\sim 45\%$ performance improvement over the next best clustering method for $K=3$. We have also conducted the same ablations over the MuJoCo setting, where the advantage of our method is also significant; see Table 4 in Appendix H.2 for more details. In addition, Figure 5 in Appendix H.2 compares a histogram of rewards for pacman and KMeans++ in the MuJoCo Halfcheetah environment, showing a consistent distributional improvement (not merely in expectation), excepting a small number of outliers.
Additionally, we consider the impact of varying budget $K$. We observe that efficacy of pacman is non-monotonic in $K$. The reason is that once $K$ is sufficiently large to cover the entire set of training tasks, increasing it further reduces the number of training tasks in individual clusters and thereby hurts generalization, both within clusters and to previously unseen tasks. Finally, we also consider the impact of changing the hyperparameter $\epsilon$, and observe that the results are relatively robust to small changes in $\epsilon$.
Overall, our method stands out by introducing only a single hyperparameter, $\epsilon$, which proves easy to tune. This simplicity is a significant advantage, especially when contrasted with the myriad of hyperparameters often encountered in typical deep RL methods. A practical approach to setting $\epsilon$ is to first compute the distances between all task embeddings and subsequently fine-tune it to ensure it remains small while still providing adequate coverage. Notably, our ablation studies further indicate that the performance of our method is far less sensitive to $\epsilon$ compared to the typical hyperparameter sensitivity observed in RL. See Appendix H for further details on these ablations.
@@ 5 Conclusion and Limitations
Acknowledgements
This work was supported by the National Science Foundation (NSF) under grants IIS-2214141 and CCF-2403758, the Office of Naval Research (ONR) under award N00014-24-1-2663, the Army Research Office (ARO) under grant W911NF-25-1-0059, and with the generous support of NVIDIA Corporation.
Impact Statement
This paper provides a theoretically grounded and empirically validated framework for personalization in Markov Decision Processes (MDPs) through the construction of policy committees.The simplicity and modularity of our approach enable wide applicability across domains such as robotics, adaptive systems, and user-centered decision-making frameworks. Furthermore, our method inherently promotes fairness by ensuring that a diverse range of tasks are represented by at least one effective policy in the committee. The clustering strategy may also extend beyond reinforcement learning to other domains where balancing efficiency and individualization is critical. By offering a scalable and provably efficient solution, our work lays a foundation for future advances in equitable and adaptive AI systems.
@@ Acknowledgements
Theorem 2.6 from (Skalse & Abate, 2023) says that two tasks $\tau_{1},\tau_{2}$ have the same ordering of policies if and only if $r_{1}$ and $r_{2}$ differ by potential shaping, positive linear scaling, and $S^{\prime}$-redistribution. Therefore, if we could find optimal policies for these two tasks separately; they necessarily differ. And by forming a policy committee of these two optimal policies, we have $V^{\Pi}_{\tau_{1}}\geq V^{\pi}_{\tau_{1}}$, $V^{\Pi}_{\tau_{2}}\geq V^{\pi}_{\tau_{2}}$, and $V^{\Pi}_{\tau_{1}}+V^{\Pi}_{\tau_{2}}>V^{\pi}_{\tau_{1}}+V^{\pi}_{\tau_{2}}$ for any policy $\pi$. ∎
@@ Appendix C Pseudocode of Greedy Intersection Algorithm
We first show $A_{1}\subset A_{2}$. Pick any $\{\theta_{1},\dots,\theta_{K}\}\in A_{1}$. Due to the premise, for each $i$, since $\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}-\epsilon\leq 0$ , there exists $\theta_{k^{*}}$ such that $\|\theta_{k^{*}}-\theta_{i}\|_{\infty}-\epsilon\leq 0$. Thus, we can have $w_{ik^{*}}=1$, and $w_{ik}=0$ for all the other $k\neq k^{*}$. Then $\mathbf{ReLU}\left(\left\{\sum_{k\in[K]}\sigma(w_{ik})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right)=\mathbf{ReLU}(\|\theta_{k^{*}}-\theta_{i}\|_{\infty}-\epsilon)=0.$ By setting $w$ this way, we could achieve the zero loss for the relaxation problem. Hence $\{\theta_{1},\dots,\theta_{K}\}\in A_{2}$.
Now to show $A_{2}\subset A_{1}$, suppose $\{\theta_{1},\dots,\theta_{K}\},w$ is a optimal solution. Due to the premise, we must have that $\mathbf{ReLU}\left(\left\{\sum_{k\in[K]}\sigma(w_{ik})\|\theta_{k}-\theta_{i}\|_{\infty}\right\}-\epsilon\right)=0$ for each $i$. Now fix $i$, since each $\sigma(w_{ik})$ is nonegative and summing them over $k$ yields $1$, there must be some positive coordinate $\sigma(w_{ik^{\prime}})$. Hence for all such $k^{\prime}$, $\mathbf{ReLU}(\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}-\epsilon)=0$, i.e., $\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}\leq\epsilon$. Thus, $\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}\leq\|\theta_{k^{\prime}}-\theta_{i}\|_{\infty}\leq\epsilon$ also holds, and $\max_{\{\theta_{1},\ldots,\theta_{K}\}}\sum_{i}\mathbf{1}(\min_{k\in[K]}\|\theta_{k}-\theta_{i}\|_{\infty}\leq\epsilon)=n$. Consequently, $\{\theta_{1},\dots,\theta_{K}\}\in A_{1}$. ∎
@@ H.2 Additional results for empirical investigation of our method
H.2.1 Clustering Ablations
We also obtained different clusters using PACMAN than using Kmeans++ and, as a result, much better performance for Halfcheetah-Velocity, as shown in the table below:
Table 4: Comparison of clustering algorithms in HalfCheetah-Velocity, Zero-Shot.
Method 6 Million Frames 12 Million Frames
KMeans++ $-105.91\pm 3.44$ $-89.50\pm 3.10$
DBScan $-234.05\pm 9.56$ $-213.42\pm 4.09$
GMM $-239.32\pm 11.27$ $-199.86\pm 9.54$
Random $-274.13\pm 16.76$ $-258.07\pm 13.62$
PACMAN -97.42 $\pm$ 3.70 -74.20 $\pm$ 6.76
As shown in Table LABEL:T:clustering_ablation_mujoco, the proposed approach outperforms all baselines.
@@ H.2.1 Clustering Ablations
The advantage of our algorithm becomes even more apparent when focusing on the welfare of the majority. To illustrate this, we present a histogram of rewards for individual test tasks during zero-shot testing using policies trained with our algorithm versus KMeans++ on the Half-Cheetah-Velocity benchmark:
@@ H.2.2 Hyperparameter Ablations
Finally, we show the effect of the $\epsilon$ hyperparameter in the Meta-World zero-shot setting. These results are reported as the success rate across all tasks for $K=2$.
@@
Table 6: Success rate for $K=2$.
[…]
500K Steps 0.23 $\pm$ 0.08 0.54 $\pm$ 0.08 0.56 $\pm$ 0.09 0.60 $\pm$ 0.07
1M Steps 0.25 $\pm$ 0.07 0.60 $\pm$ 0.10 0.58 $\pm$ 0.14 0.58 $\pm$ 0.07
Table 7: Success rate for $K=3$.
@@ Appendix I Meta-World Task Descriptions
When $K=3$, the three Meta-World clusters (for one of the random subsamples of 30 tasks) are provided below.
Cluster 1 Cluster 2 Cluster 3
handle-press-side-v1 peg-unplug-side-v1 pick-out-of-hole-v1 pick-place-v1 plate-slide-back-v1 plate-slide-side-v1 plate-slide-v1 push-back-v1 push-v1 reach-v1 stick-push-v1 faucet-open-v1 hammer-v1 handle-press-side-v1 handle-press-v1 handle-pull-side-v1 peg-insert-side-v1 peg-unplug-side-v1 pick-out-of-hole-v1 pick-place-v1 pick-place-wall-v1 plate-slide-back-side-v1 plate-slide-side-v1 plate-slide-v1 soccer-v1 sweep-v1 hand-insert-v1 handle-pull-v1 lever-pull-v1 peg-insert-side-v1 push-wall-v1 reach-wall-v1 shelf-place-v1 stick-pull-v1 stick-push-v1 sweep-into-v1 window-close-v1 window-open-v1
Table 9: Meta-World Task Clusters
On the high level, the clustering can be described as corresponding to three categories of manipulation tasks. The first cluster involves pulling and sliding manipulations (e.g., pull a puck, move end-effector, slide a plate to and from the cabinet). The second cluster includes manipulations that involve pressing and pushing (e.g., rotate the faucet, press a handle down, insert a peg into a hole, kick a soccer ball, sweep a puck). The third cluster consists of tasks involving indirect or constrained manipulation and which thus require better precision and control (e.g., bypass a wall, grasp a stick and pull or push a box with it).